# Multispectral airborne laser scanning for tree species classification: a benchmark of machine learning and deep learning algorithms

### v1

@@ 4.2.2 Scaling laws with respect to point density
We also study the classification error as a function of point density by randomly sub-sampling the segments to a desired point density as illustrated in Fig. 11(a). We vary the point density from $10^{-1}$ $\mathrm{pts}/\mathrm{m}^{2}$ representing an extremely sparse point cloud to around 1000 $\mathrm{pts}/\mathrm{m}^{2}$ representing a dense ALS point cloud. Here, we evaluate the classification performance using 5-fold cross-validation based on the full benchmarking dataset similarly to Sec. 4.2.1.
@@
In Figs. 11(b)-(d), we compare the classification error as a function of point density for FGI-PointTransformer-DL-3D representing a baseline DL model, FGI-PointTransformer-GeometryOnly-DL-3D using only geometric information, a variant of the baseline DL model using only 25% of the training set, and FGI-RF-ML representing a ML model. We observe that the baseline DL model exhibits two regimes of convergence. At low point densities below 10 $\mathrm{pts}/\mathrm{m}^{2}$, the error of the baseline DL model exhibits rapid convergence following a power law
@@
Importantly, the geometry-only DL model exhibits 2–4 times higher overall and macro-average classification error across the studied point densities compared to the baseline DL model utilizing multi-spectral information as shown in Fig. 11(b). Interestingly, the geometry-only DL model shows a slower convergence of the classification error at low point densities compared to the baseline DL model, but the power law scaling extends across a larger range of point densities up to $\geq 100$ $\mathrm{pts}/\mathrm{m}^{2}$. Thus, the geometry-only DL model begins to catch up with the baseline model for point densities above 10 $\mathrm{pts}/\mathrm{m}^{2}$. This also suggests that the improvement provided by the multi-spectral information is the largest for point densities around 10 $\mathrm{pts}/\mathrm{m}^{2}$, which highlights the advantage of multi-spectral information especially on sparse ALS datasets. For the largest considered point density of around 500 $\mathrm{pts}/\mathrm{m}^{2}$, the overall classification error of the geometry-only DL model still remains a factor of two higher than that of the baseline DL model using multispectral information. The observed results confirm similar findings from a multitude of applications relating to multispectral- and hyperspectral laser scanning (Takhtkeshha et al., 2024).
@@ 4.3 Effect of segmentation quality on classification accuracy
In this section, we study the classification accuracy across different profile categories and tree crown classes introduced in Sec. 2.6, see also Fig. 5. We show the overall classification accuracy for the best deep learning method and the best machine learning method in the benchmarking study together with the corresponding median values across the five profile categories using the HeliALS data in Fig. 12(a) and using the Optech Titan data in Fig. 12(b). For both the HeliALS and Optech Titan datasets, the highest overall accuracy is reached for the categories ‘Single tree‘ and ‘Large tree with undergrowth‘, which cover a large majority of the dataset as shown in Fig. 5(b). This is expected since these profile categories represent ideal segments with no nearby growing trees of the same or different species. On both datasets, the worst classification accuracy is obtained for the profile category ‘Many trees‘ representing a segment, in which the largest tree is accompanied with one or multiple smaller trees of different species. Nevertheless, the best deep learning model using the HeliALS dataset is able to predict the species of the largest tree correctly in such segments with an overall (macro-average) accuracy of 75.5% (64.9%).
@@
For the HeliALS dataset, we observe that the best DL method consistently outperforms the best ML method across the profile categories as shown in Fig. 12(a). When considering the macro-average accuracy (see A.), the improvement provided by the DL method becomes more apparent especially for the most challenging category ‘Many trees‘. For this category, the best DL method reaches a macro-average accuracy of 64.9%, while the corresponding quantity for the best ML method is only 45.5%. For the Optech Titan dataset, we observe that the best DL and ML methods perform equivalently across the different profile categories as shown in Fig. 12(b), though the DL model provides a small improvement for the category ‘Many trees‘, while the roles are reversed for the category ‘Tree section‘.
@@
When it comes to the crown class categories, we observe that the overall classification accuracy is the highest and essentially equivalent for the classes ‘Isolated‘, ‘Dominant‘, and ‘Roadside trees‘ across the different deep learning and machine learning models on both datasets as shown in Figs. 13(a) and (b). For the crown class ‘Co-dominant‘, the overall accuracy is a few percentage points lower compared to the above mentioned crown classes. Importantly, the classification accuracy is significantly lower for the category ‘Smaller Tree next to larger tree‘. For example, the best DL method on the HeliALS data reaches an overall accuracy of only 67% for this category, which is over 20 percentage points lower than the overall accuracy of around 90% for the crown classes ‘Isolated‘, ‘Dominant‘, and ‘Roadside trees‘. These observations are consistent with prior research (e.g., Hakula et al., 2023), where suppressed trees, characterized by their small size and proximity to larger neighbors, had an overall accuracy 20 percentage points lower than that of dominant trees.
@@
To the best of our knowledge, the effect of segmentation quality on species classification has not been investigated in detail in prior literature since the segmentation has often been improved manually to ensure near-ideal tree segments, such as in Puliti et al. (2025). However, ALS-based tree species estimation across large areas will necessitate automatic segmentation, which will unavoidably lead to imperfect segmentation quality, especially for small or suppressed trees. This highlights the relevance of studying the effect of segmentation quality and crown class on species classification. In our study, the segmentation was based on a traditional watershed approach applied to the sparse Optech Titan dataset. In the future, segmentation methods based on deep learning, such as ForAINet (Xiang et al., 2024), SegmentAnyTree (Wielgosz et al., 2024), and TreeLearn (Henrich et al., 2024), may improve the segmentation quality especially on dense ALS data, where the high point density may contain cues for improved segmentation beyond that achievable with traditional approaches.
@@ 4.4 Further Discussion
Our study compares favorably against recent UAV-based studies employing state-of-the-art laser scanning systems and high-quality passive hyperspectral imaging systems. For example, Quan et al. (2023) used a Random Forest classifier with leaf-on data acquired from a Riegl VUX-1UAV laser scanner and a MicroCASI 1920 hyperspectral imager to classify 11 species (2 coniferous, 9 broad-leaved), achieving a macro-average accuracy of 69.2% on a dataset with 904 trees using 90/10 train-test split. Zhong et al. (2022) employed a Support Vector Machine to classify 5 species (1 coniferous, 4 broad-leaved) using data from a RIEGL miniVUX-1UAV and a Resonon Pika L hyperspectral sensor. They reported an overall accuracy of 89.2% on a dataset with 880 trees using a 60/40 train-test split. In our study, the best-performing model, FGI-PointTransformer-DL, reaches an overall accuracy of 92.0% and a macro-average accuracy of 85.1% across the nine studied species when using the full HeliALS dataset with 6256 trees and 80/20 train-test split, see Fig. 9(a). Thus, our best result surpasses the performance of these recent state-of-the-art studies on UAV-based species classification.
@@
To gain further insights into the classification errors discussed in Sec. 4.1, we visualize the clustering of different tree species in the feature space in Figs. 14(a) and (b) using input features of FGI-RF-ML and the deep features of FGI-PointTransformer-DL-3D, respectively. To project the high-dimensional features into a two-dimensional plane, we use t-distributed stochastic neighbor embedding (t-SNE) (Hinton and Roweis, 2002) with perplexity values ranging from 15 to 30 to have a clear separation of the main clusters. In Fig. 14(a) showing the clusters for the RF features, we observe that pines, spruces, birches and maples mostly form separate clusters in line with the high classification accuracy of these species. Visually, these species have clearly distinct characteristics, and they are easy to recognize in the field based on their shapes and needles or leaves, see Fig. 3. However, we note that a small set of spruces is located in the middle of the feature space between pine and alder clusters, separate from other spruces. A closer inspection of individual trees in this cluster showed that these spruces were typically growing on yards, roadsides or other places in the built-up environment, and they had a narrow shape different from large spruces growing in a forest.
For the remaining five species, there is significant inter-species overlap in the feature space. Clusters for aspen and alder are located next to each other with a partial overlap, which seems natural due to the similar appearance of these species. Furthermore, aspens also have overlap with birches in the feature space. Rowans, oaks and lindens are overlapping and close to other deciduous species, especially maples. This might be related to similarities in the size and shape of these species. Naturally, these observations are well-aligned with our remarks based on the confusion matrices presented in Sec. 4.1. Importantly, the separation of deciduous species in the feature space and hence their classification accuracy may be affected by the season of data acquisition. For example, Lisiewicz et al. (2025) observed an increase in the classification accuracy of oak and aspen using ALS data collected during leaf-off season, whereas maple and linden were more accurately classified using leaf-on ALS data. A combination of features derived from leaf-off and leaf-on data has been shown to improve the species classification accuracy in several studies (Kim et al., 2009; Shi et al., 2018b; Kamińska et al., 2021; Lisiewicz et al., 2025).
@@
As briefly discussed in Sec. 4.1, DetailView-DL-2D trained on our dataset without any spectral information achieved a comparable classification accuracy as the Point Transformer model augmented with single-channel (VUX) reflectance. Hence, further studies should investigate incorporating single-channel or multi-channel spectral information in the DetailView model, which holds potential to improve the classification accuracy especially for the minority species. Importantly, retraining of the DetailView model was necessary to achieve competitive performance on the ALS data of the current study. This conclusion seems to be valid even when accounting for the different species composition between our study and that of Puliti et al. (2025), which we tested by re-training the classification head while using frozen encoder parameters of the sensor-agnostic model from Puliti et al. (2025). Such an approach resulted in an overall (macro-average) accuracy of 78.7% (56.7%), which is still significantly lower compared to the overall (macro-average) accuracy of 84.3% (63.9%) for the fully retrained model.
Based on our studies with varying point densities in Sec. 4.2.2, the influence of single- and multi-channel spectral information is most pronounced at low point densities of around $10~{}\mathrm{pts}/\mathrm{m}^{2}$. However, we still observe a significant improvement in the classification accuracy when using the HeliALS data ($>1000$ pts/$\mathrm{m}^{2}$) instead of the Optech Titan data (35 pts/$\mathrm{m}^{2}$), especially when it comes to the macro-average accuracy that is 14 percentage points higher for the best-performing model of the HeliALS data. Importantly, our studies with a varying point density rely on artificial subsampling of the HeliALS point cloud. In practical implementations, the reduced point density is often a byproduct of a higher flight altitude, which may reduce the point cloud quality due to an increased beam width. Thus, our results in Fig. 11 may not perfectly describe a scenario, in which the flight altitude controls the point density. It is also important to note the disparity in the availability of point cloud data with densities similar to HeliALS or Optech Titan systems. The Optech Titan data exhibits a density comparable to that used in municipal mapping applications in Finland, typically ranging from 40 to 60 points/$\mathrm{m}^{2}$. Conversely, the HeliALS data density aligns more closely with drone-based LiDAR surveys, indicating its suitability for smaller, more localized studies, e.g. for forest holding.
In principle, our results demonstrate precision and recall values suitable for operational mapping of aspen populations at the individual tree level, reaching $F_{1}=0.71$ for the best performing model on the HeliALS data and $F_{1}=0.57$ for the best-performing model on the Optech Titan data. However, the proportion of aspen trees was higher on our test site compared to typical boreal forests, and therefore, a lower $F_{1}$ score may be achieved in operational use for an equivalent training set size. For example, Toivonen et al. (2024) reached an $F_{1}$ score of 0.44 for mapping large aspens using ALS (7.2 $\mathrm{pts}/\mathrm{m}^{2}$) and aerial imagery on sample plots with an aspen proportion of around 6%, whereas the $F_{1}$ score reduced to 0.21 on sample plots with an aspen proportion of around 0.1%. Nevertheless, our results suggest that a higher point density and multispectral information improve the classification accuracy of aspen.
@@
Multispectral data also opens up new possibilities for the use of echo classes in species classification. After the first echo, the LiDAR signal is affected by a variety of factors, such as transmission, absorption losses, and multiple scattering, which alter the measured return intensity. Importantly, the return intensities in such multiple scattering events are affected differently for each wavelength, and the effect may vary by species due to different canopy geometries, thus providing additional information for species classification. For example, features related to intermediate echos were found to contribute to the classification performance of UEF-RF-ML, though features related to the first echoes were the most important ones as they correspond to the purest return intensities. Features related to last echos provided no added value to the classifier.
More studies are also needed to investigate the transferability of deep learning methods to new datasets that may also cover large areas. In Sec. 4.2.1, we observed that the classification error followed a power law scaling as a function of training set size, which, in principle, enables us to estimate the required number of training trees for a desired classification accuracy when using a given model. However, we also noted that the DetailView model performed significantly worse in this study when trained on the FOR-species20K dataset compared to the training set of the current study. Thus, it remains to be verified whether a similar power law scaling of the classification error would be observed if the models trained on the current dataset were transferred to new areas, though in the same forest zone. The gap between deep learning and machine learning models may narrow down if the faster convergence of deep learning models is partially due to overfitting to the current dataset.

### v2

@@ 4.4.2 Scaling laws with respect to point density
We also study the classification error as a function of point density by randomly sub-sampling the segments to a desired point density as illustrated in Fig. 11(a) and explained in Sec. 3.4. In Figs. 11(b)-(d), we compare the classification error as a function of point density for FGI-PointTransformer-DL-3D representing a baseline DL model, FGI-PointTransformer-GeometryOnly-DL-3D using only geometric information, a variant of the baseline DL model using only 25% of the training set, and FGI-RF-ML representing a ML model. We observe that the baseline DL model exhibits two regimes of convergence. At low point densities below 10 $\mathrm{pts}/\mathrm{m}^{2}$, the error of the baseline DL model exhibits rapid convergence following a power law
@@
The geometry-only DL model exhibits 2–4 times higher overall and macro-average classification error across the studied point densities compared to the baseline DL model utilizing multi-spectral information as shown in Fig. 11(b). Interestingly, the geometry-only DL model shows a slower convergence of the classification error at low point densities compared to the baseline DL model, but the power law scaling extends across a larger range of point densities up to $\geq 100$ $\mathrm{pts}/\mathrm{m}^{2}$. Thus, the geometry-only DL model begins to catch up with the baseline model for point densities above 10 $\mathrm{pts}/\mathrm{m}^{2}$. This also suggests that the improvement provided by the multi-spectral information is the largest for point densities around 10 $\mathrm{pts}/\mathrm{m}^{2}$, which highlights the advantage of multi-spectral information especially on sparse ALS datasets. For the largest considered point density of around 500 $\mathrm{pts}/\mathrm{m}^{2}$, the overall classification error of the geometry-only DL model still remains a factor of two higher than that of the baseline DL model using multispectral information. The observed results confirm similar findings from a multitude of applications relating to multispectral- and hyperspectral laser scanning (takhtkeshha2024multispectral).
@@ 4.5 Effect of segmentation quality on classification accuracy
In this section, we study the classification accuracy across different profile categories and tree crown classes introduced in Sec. 2.6, see also Fig. 5. We show the overall classification accuracy for the best deep learning method and the best machine learning method in the benchmarking study together with the corresponding median values across the five profile categories using the HeliALS data in Fig. 12(a) and using the Optech Titan data in Fig. 12(b). For both the HeliALS and Optech Titan datasets, the highest overall accuracy is reached for the categories ‘Single tree‘ and ‘Large tree with undergrowth‘, which cover a large majority of the dataset as shown in Fig. 5(b). This is expected since these profile categories represent ideal segments with no nearby growing trees of the same or different species. On both datasets, the worst classification accuracy is obtained for the profile category ‘Many trees of many species‘ representing a segment, in which the largest tree is accompanied with one or multiple smaller trees of different species. Nevertheless, the best deep learning model using the HeliALS dataset is able to predict the species of the largest tree correctly in such segments with an overall (macro-average) accuracy of 75.5% (64.9%).
@@
For the HeliALS dataset, we observe that the best DL method consistently outperforms the best ML method across the profile categories as shown in Fig. 12(a). When considering the macro-average accuracy (see C), the improvement provided by the DL method becomes more apparent especially for the most challenging category ‘Many trees of many species‘. For this category, the best DL method reaches a macro-average accuracy of 64.9%, while the corresponding quantity for the best ML method is only 45.5%. For the Optech Titan dataset, we observe that the best DL and ML methods perform equivalently across the different profile categories as shown in Fig. 12(b), though the DL model provides a small improvement for the category ‘Many trees of many species‘, while the roles are reversed for the category ‘Tree section‘.
@@
When it comes to the crown class categories, we observe that the overall classification accuracy is the highest and essentially equivalent for the classes ‘Isolated‘, ‘Dominant‘, and ‘Roadside trees‘ across the different deep learning and machine learning models on both datasets as shown in Figs. 13(a) and (b). For the crown class ‘Co-dominant‘, the overall accuracy is a few percentage points lower compared to the above mentioned crown classes. The classification accuracy is significantly lower for the category ‘Smaller Tree next to larger tree‘. For example, the best DL method on the HeliALS data reaches an overall accuracy of only 67% for this category, which is over 20 percentage points lower than the overall accuracy of around 90% for the crown classes ‘Isolated‘, ‘Dominant‘, and ‘Roadside trees‘. These observations are consistent with prior research (e.g., hakula2023individual), where suppressed trees, characterized by their small size and proximity to larger neighbors, had an overall accuracy 20 percentage points lower than that of dominant trees.
@@
To the best of our knowledge, the effect of segmentation quality on species classification has not been investigated in detail in prior literature since the segmentation has often been improved manually to ensure near-ideal tree segments, such as in puliti2025benchmarking. However, ALS-based tree species estimation across large areas will necessitate automatic segmentation, which will unavoidably lead to imperfect segmentation quality, especially for small or suppressed trees. In our study, the segmentation was based on a traditional watershed approach applied to the sparse Optech Titan dataset. In the future, segmentation methods based on deep learning, such as ForAINet (xiang2024automated), SegmentAnyTree (wielgosz2024segmentanytree), and TreeLearn (henrich2024treelearn), may improve the segmentation quality especially on dense ALS data, where the high point density may contain cues for improved segmentation beyond traditional approaches. Further work should also study whether it is advisable to use the same segmentation method for both the training phase and the prediction phase in an operational setting as opposed to using a high-quality, labour-intensive manual segmentation for the training phase and an automatic segmentation for the prediction phase. According to our hypothesis, it is better to use the same segmentation method throughout the process to ensure that the training and prediction data come from the same distribution.
@@ 4.6 Further Discussion
Our study compares favorably against recent UAV-based studies employing state-of-the-art laser scanning systems and high-quality passive hyperspectral imaging systems. For example, quan2023tree used a Random Forest classifier with leaf-on data acquired from a Riegl VUX-1UAV laser scanner and a MicroCASI 1920 hyperspectral imager to classify 11 species (2 coniferous, 9 broad-leaved), achieving a macro-average accuracy of 69.2% on a dataset with 904 trees using 90/10 train-test split. zhong2022identification employed a Support Vector Machine to classify 5 species (1 coniferous, 4 broad-leaved) using data from a RIEGL miniVUX-1UAV and a Resonon Pika L hyperspectral sensor. They achieved an overall accuracy of 89.2% and a macro-average accuracy of 84.5% on a dataset with 880 trees using a 60/40 train-test split. In our study, the best-performing model, FGI-PointTransformer-DL, reaches an overall accuracy of 92.0% and a macro-average accuracy of 85.1% across the nine studied species when using the full HeliALS dataset with 6257 segments and 80/20 train-test split, see Fig. 9(a). Thus, our best result surpasses the performance of these recent state-of-the-art studies on UAV-based species classification.
@@
To gain further insights into the classification errors discussed in Sec. 4.3, we visualize the clustering of different tree species in the feature space in Figs. 14(a) and (b) using input features of FGI-RF-ML and the deep features of FGI-PointTransformer-DL-3D, respectively. To project the high-dimensional features into a two-dimensional plane, we use t-distributed stochastic neighbor embedding (t-SNE) (hinton2002stochastic) with perplexity values ranging from 15 to 30 to have a clear separation of the main clusters. In Fig. 14(a) showing the clusters for the RF features, we observe that pines, spruces, birches and maples mostly form separate clusters in line with the high classification accuracy of these species. Visually, these species have clearly distinct characteristics, and they are easy to recognize in the field based on their shapes and needles or leaves, see Fig. 3. However, we note that a small set of spruces is located in the middle of the feature space between pine and alder clusters, separate from other spruces. A closer inspection of individual trees in this cluster showed that these spruces were typically growing on yards, roadsides or other places in the built-up environment, and they had a narrow shape different from large spruces growing in a forest.
For the remaining five species, there is significant inter-species overlap in the feature space. Clusters for aspen and alder are located next to each other with a partial overlap, which seems natural due to the similar appearance of these species. Furthermore, aspens also have overlap with birches in the feature space. Rowans, oaks and lindens are overlapping and close to other deciduous species, especially maples. This might be related to similarities in the size and shape of these species. Naturally, these observations are well-aligned with our remarks based on the confusion matrices presented in Sec. 4.3. The separation of deciduous species in the feature space and hence their classification accuracy may be affected by the season of data acquisition. For example, lisiewicz2025comprehensive observed an increase in the classification accuracy of oak and aspen using ALS data collected during leaf-off season, whereas maple and linden were more accurately classified using leaf-on ALS data. A combination of features derived from leaf-off and leaf-on data has been shown to improve the species classification accuracy in several studies (kim2009tree; shi2018important; kaminska2021single; lisiewicz2025comprehensive).
@@
As briefly discussed in Sec. 4.1, DetailView-DL-2D trained on our dataset without any spectral information achieved a comparable classification accuracy as the Point Transformer model augmented with single-channel (VUX) reflectance. Hence, further studies should investigate incorporating single-channel or multi-channel spectral information in the DetailView model, which holds potential to improve the classification accuracy especially for the minority species. Importantly, retraining of the DetailView model was necessary to achieve competitive performance on the ALS data of the current study. This conclusion seems to be valid even when accounting for the different species composition between our study and that of puliti2025benchmarking, which we tested by re-training the classification head while using frozen encoder parameters of the sensor-agnostic model from puliti2025benchmarking. Such an approach resulted in an overall (macro-average) accuracy of 78.7% (56.7%), which is still significantly lower compared to the overall (macro-average) accuracy of 84.3% (63.9%) for the fully retrained model.
[…]
In principle, our results demonstrate precision and recall values suitable for operational mapping of aspen populations at the individual tree level, reaching $F_{1}=0.71$ for the best performing model on the HeliALS data and $F_{1}=0.57$ for the best-performing model on the Optech Titan data. However, the proportion of aspen trees was higher on our study area compared to typical boreal forests, and therefore, a lower $F_{1}$ score may be achieved in operational use for an equivalent training set size. For example, toivonen2024mapping reached an $F_{1}$ score of 0.44 for mapping large aspens using ALS (7.2 $\mathrm{pts}/\mathrm{m}^{2}$) and aerial imagery on sample plots with an aspen proportion of around 6%, whereas the $F_{1}$ score reduced to 0.21 on sample plots with an aspen proportion of around 0.1%. Nevertheless, our results suggest that a higher point density and multispectral information improve the classification accuracy of aspen.
@@
When it comes to mapping large areas, the computational cost of species classification becomes also a relevant performance metric even though species classification for forest inventory is typically a post-processing activity without strict runtime requirements. To assess the scalability of the studied classification methods, we compare the run time of model training and inference in B for three selected methods including FGI-RF-ML representing a shallow machine learning method, FGI-PointTransformer-DL-3D representing a 3D deep learning method, and DetailView-DL-2D representing a 2D deep learning method. We found that the training time of the RF model was the shortest, but interestingly, the inference time of FGI-PointTransformer-DL-3D was 177 times shorter than that of the RF model and 52 times shorter than that of the DetailView on the HeliALS data. For the RF model, the feature computation dominated the total runtime, whereas the training and inference times after the feature computation were quick. For more details, see Appendix B and Fig. B.1.
Multispectral data also opens up new possibilities for the use of echo classes in species classification. After the first echo, the LiDAR signal is affected by a variety of factors, such as transmission, absorption losses, and multiple scattering, which alter the measured return intensity. The return intensities in such multiple scattering events are affected differently for each wavelength, and the effect may vary by species due to different canopy geometries, thus providing additional information for species classification. For example, features related to intermediate echos were found to contribute to the classification performance of UEF-RF-ML, though features related to the first echoes were the most important ones as they correspond to the purest return intensities. Features related to last echos provided no added value to the classifier.
More studies are also needed to investigate the transferability of deep learning methods to new datasets that may also cover large areas. In Sec. 4.4.1, we observed that the classification error followed a power law scaling as a function of training set size, which, in principle, enables us to estimate the required number of training segments for a desired classification accuracy when using a given model. However, we also noted that the DetailView model performed significantly worse in this study when trained on the FOR-species20K dataset compared to the training set of the current study. Thus, it remains to be verified whether a similar power law scaling of the classification error would be observed if the models trained on the current dataset were transferred to new areas, though in the same forest zone. The gap between deep learning and machine learning models may narrow down if the faster convergence of deep learning models is partially due to overfitting to the current dataset.